### v1

# Engineering and Computer Science, DGIST, South Korea Bridging Geometric and Semantic Foundation Models

Abstract
We present Bridging Geometric and Semantic (BriGeS), an effective method that fuses geometric and semantic information within foundation models to enhance Monocular Depth Estimation (MDE). Central to BriGeS is the Bridging Gate, which integrates the complementary strengths of depth and segmentation foundation models. This integration is further refined by our Attention Temperature Scaling technique. It finely adjusts the focus of the attention mechanisms to prevent over-concentration on specific features, thus ensuring balanced performance across diverse inputs. BriGeS capitalizes on pre-trained foundation models and adopts a strategy that focuses on training only the Bridging Gate. This method significantly reduces resource demands and training time while maintaining the model’s ability to generalize effectively. Extensive experiments across multiple challenging datasets demonstrate that BriGeS outperforms state-of-the-art methods in MDE for complex scenes, effectively handling intricate structures and overlapping objects.
1 Introduction
Figure 1: Qualitative comparison with DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao]. Our method excels in capturing more complex visual elements and details.
Monocular Depth Estimation (MDE) is a crucial task that predicts the depth map from a single image. This method is cost-effective and simple to deploy. Utilizing just one camera eliminates the need for complex hardware setups or additional camera installations, thereby reducing initial expenses and simplifying maintenance. Given its simple structure, it can be seamlessly integrated across various environments and platforms. This adaptability makes it especially valuable in fields like robotics, autonomous driving, and augmented reality.
Historically, MDE relies on handcrafted features and traditional vision techniques [Hoiem et al.(2007)Hoiem, Efros, and Hebert, Liu et al.(2008)Liu, Yuen, Torralba, Sivic, and Freeman, Saxena et al.(2008)Saxena, Sun, and Ng]. However, the advent of CNN [Eigen et al.(2014)Eigen, Puhrsch, and Fergus, Laina et al.(2016)Laina, Rupprecht, Belagiannis, Tombari, and Navab, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Wang et al.(2020b)Wang, Zhang, Wang, Lu, and Ruan, Ricci et al.(2018)Ricci, Ouyang, Wang, Sebe, et al., Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow] and ViT-based models [Yang et al.(2021)Yang, Tang, Ding, Sebe, and Ricci, Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun, Yuan et al.(2022)Yuan, Gu, Dai, Zhu, and Tan, Zhao et al.(2022)Zhao, Zhang, Poggi, Tosi, Guo, Zhu, Huang, Tang, and Mattoccia, Cheng et al.(2021)Cheng, Zhang, and Tang, Bae et al.(2023)Bae, Moon, and Im] has significantly enhanced its performance. Recent studies [Kendall et al.(2017)Kendall, Martirosyan, Dasgupta, Henry, Kennedy, Bachrach, and Bry, Klingner et al.(2020)Klingner, Termöhlen, Mikolajczyk, and Fingscheidt, Wang et al.(2020a)Wang, Zhang, Wang, Lin, and Lu, Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li, Xing et al.(2023)Xing, Shen, Ho, and Tzes, Shvets et al.(2024)Shvets, Zhao, Niethammer, Sengupta, and Berg, Li et al.(2023)Li, Xue, Su, He, Mao, Zhu, Sun, and Zhang, Choi et al.(2020)Choi, Jung, Lee, and Kim, Sanchez-Escobedo et al.(2018)Sanchez-Escobedo, Lin, Casas, and Pardas, Zama Ramirez et al.(2019)Zama Ramirez, Poggi, Tosi, Mattoccia, and Di Stefano, Jiao et al.(2018)Jiao, Cao, Song, and Lau, Chen et al.(2019)Chen, Liu, Liu, and Wang] have highlighted the advantages of incorporating semantic context from images, which provide more informative cues than solely relying on local geometry. More recently, MDE works [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have focused on training large-scale models using a vast amount of training datasets inspired by foundation models [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al., Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun, Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]. These approaches aim to harness the broader contextual understanding inherent in large pre-trained models, significantly advancing the field of depth estimation.
Despite these advancements, current foundation models in MDE [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have primarily utilized geometric data for depth estimation and have not directly incorporated semantic information. These models require substantial training data and considerable computational resources, making it challenging to integrate additional semantic information. In response to these challenges, we propose a new integration module for MDE that adaptively incorporates semantic information, named Bridging Geometric and Semantic (BriGeS). This approach is particularly effective in addressing complex visual elements such as intricate structures, homogeneous regions, and delicate structures as shown in Fig. 1.
To integrate depth features with semantic features under a resource-constrained training environment, we design Bridging Gate, which directly fuses features from the depth [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] and segmentation [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] foundation model. This gate employs cross-attention and self-attention blocks to facilitate effective feature integration. To ensure high-depth estimation performance while preserving prior knowledge of a model trained on large-scale data, we exclusively train the Bridging Gate, keeping all encoders and decoders frozen. Additionally, we introduce the Attention Temperature Scaling technique to adjust the variance of the Bridging Gate’s attention map. This adjustment is crucial as the map tends to become overly concentrated on specific regions due to the fusion of two distinct modalities. By altering the distribution of attention operations during the inference stage, the technique mitigates this issue, thereby simplifying the process and enhancing overall effectiveness.
Extensive experiments demonstrate significant performance improvements achieved by employing BriGeS on unseen datasets, particularly excelling in images with complex structures. Our contributions to the field can be summarized as follows:
  • We present BriGeS, an effective module that fuses a depth foundation model and a segmentation foundation model using minimal data and training effort to enhance Monocular Depth Estimation.
  • We propose the Bridging Gate, an adaptive fusion layer specifically tailored to integrate geometric and semantic information.
  • We introduce the Attention Temperature Scaling, which regulates the attention distribution to mitigate over-concentration on specific regions during inference.
2 Related Works
Monocular Depth Estimation. Monocular Depth Estimation (MDE) method estimates depth from a single image by learning the spatial shape and structure of objects. Some pioneering research in this field predominantly utilizes handcrafted features and traditional vision techniques [Hoiem et al.(2007)Hoiem, Efros, and Hebert, Liu et al.(2008)Liu, Yuen, Torralba, Sivic, and Freeman, Saxena et al.(2008)Saxena, Sun, and Ng]. However, these methods often rely solely on local information and face challenges in capturing comprehensive scene structure. To address these challenges, many MDE models have adopted CNN [Eigen et al.(2014)Eigen, Puhrsch, and Fergus, Laina et al.(2016)Laina, Rupprecht, Belagiannis, Tombari, and Navab, Ricci et al.(2018)Ricci, Ouyang, Wang, Sebe, et al., Wang et al.(2020b)Wang, Zhang, Wang, Lu, and Ruan] and ViT-based approaches [Yang et al.(2021)Yang, Tang, Ding, Sebe, and Ricci, Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun, Yuan et al.(2022)Yuan, Gu, Dai, Zhu, and Tan, Bae et al.(2023)Bae, Moon, and Im]. Recently, research has pivoted towards a self-supervised approach [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Zhou et al.(2022)Zhou, Taylor, Greenwood, and Mackiewicz, Fan et al.(2023)Fan, Tosi, Poggi, and Mattoccia, Saunders et al.(2024)Saunders, Manso, and Vogiatzis] due to the high cost of labeling training data, allowing training on larger datasets. While these models face challenges with objects having unclear boundaries or similar depths [Klingner et al.(2020)Klingner, Termöhlen, Mikolajczyk, and Fingscheidt, Dinh and Lim(2023), Shvets et al.(2024)Shvets, Zhao, Niethammer, Sengupta, and Berg], numerous studies have started integrating semantic information to refine edge predictions [Wang et al.(2020a)Wang, Zhang, Wang, Lin, and Lu, Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li, Xing et al.(2023)Xing, Shen, Ho, and Tzes]. For instance, ROIFormer [Xing et al.(2023)Xing, Shen, Ho, and Tzes] proposes a semantic-aware region of interest transformer that uses geometric cues from semantic segmentation to guide the attention mechanism. TriDepth [Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li] further introduces a semantic-aware triplet loss to address the edge-fattening problem, which significantly improves model performance while incurring no additional computational cost.
Zero-shot Depth Estimation. Zero-shot depth estimation involves predicting depth from images unseen during training. Earlier efforts in this domain focused on collecting more training images to enhance the generalization capabilities of MDE models [Chen et al.(2016)Chen, Fu, Yang, and Deng, Xian et al.(2018)Xian, Shen, Cao, Lu, Xiao, Li, and Luo, Chen et al.(2020)Chen, Qian, Fan, Kojima, Hamilton, and Deng]. However, data variability in unconstrained settings often leads to inconsistent depth prediction due to differing depth scales and shifts. To resolve these problems, several methods are designed to estimate affine-invariant depth, which adjusts for global scale and shift discrepancies without knowing the precise object size. A pioneering work, MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] proposes an affine-invariant loss to facilitate effective zero-shot transfer across diverse datasets by estimating relative depth. DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun] also adapts the same loss term, and further enhances the performance of zero-shot MDE by replacing CNN-based architecture with vision transformers [Dosovitskiy(2020)]. For metric depth estimation, ZoeDepth [Bhat et al.(2023)Bhat, Birkl, Wofk, Wonka, and Müller] integrates relative and metric depth supervision, and Metric3D [Yin et al.(2023)Yin, Zhang, Chen, Cai, Yu, Wang, Chen, and Shen, Hu et al.(2024)Hu, Yin, Zhang, Cai, Long, Chen, Wang, Yu, Shen, and Shen] introduces a canonical camera space transformation module to resolve metric ambiguity. Recent works leverage the rich visual knowledge from the pre-trained model to perform zero-shot knowledge transfer. For instance, DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] develops a semi-supervised training strategy using DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] encoder, while Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] and GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen] use a model derived from Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer].
3 Preliminary: Foundation Models
In this section, we provide a brief overview of the foundation model used in our pipeline.
DepthAnything. The depth foundation model, DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] employs a ViT-based encoder $E_{d}$ and a CNN-based decoder $D_{d}$. The encoder processes the input image $I$ to generate features $\{f^{i}_{d}\}_{i=1}^{4}$, which serve as the input for the decoder. The decoder then utilizes this feature to regress the depth map $I_{d}$ of the input image as follows:
$\begin{gathered}I_{d}=D_{d}(\{f^{i}_{d}\}_{i=1}^{4}),~{}\{f^{i}_{d}\}_{i=1}^{4}=E_{d}(I),\\ I_{d}\in\mathbb{R}^{H^{\prime}\times W^{\prime}},~{}f^{i}_{d}\in\mathbb{R}^{HW\times C},~{}I\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3}\end{gathered}$ (1)
where $H$, $W$, and $C$ denote the height, width, and channel of the feature map, respectively, while $H^{\prime}$ and $W^{\prime}$ represent the height and width of the input image and the resulting depth.
SegmentAnything. The segmentation foundation model, SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] consists of a ViT-based encoder $E_{s}$, a prompt encoder $E_{p}$, and a decoder $D_{s}$. The encoder $E_{s}$ processes the input image $I$ to generate a feature $f_{s}$. The prompt encoder takes an input $x$, which can be points, boxes, or text, and generates a corresponding prompt token $p$. Finally, the decoder combines the feature map $f_{s}$ with the prompt token $p$ to generate the mask $M$ for the input image as follows:
$\begin{gathered}M=D_{s}(f_{s},p),~{}f_{s}=E_{s}(I),~{}p=E_{p}(x),\\ ~{}M\in\mathbb{R}^{H^{\prime}\times W^{\prime}},~{}f_{s}\in\mathbb{R}^{C\times H\times W},I\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3},\end{gathered}$ (2)
where $H^{\prime}$ and $W^{\prime}$ represent the height and width of masks.
4 Method
4.1 Overall Pipeline
In this section, we describe the overall pipeline of BriGeS, illustrated in Fig. 2. Our BriGeS utilizes the encoders $E_{d}$ from DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] and $E_{s}$ from SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.], along with the DepthAnything decoder $D_{d}$. We extract depth features $\{f^{i}_{d}\}_{i=1}^{4}$ and semantic feature $f_{s}$ from the respective encoders, and fuse them using Bridging Gates $\{G^{i}\}_{i=1}^{4}$ to generate semantic-aware geometric features $\mathbf{F}_{sg}=\{F^{i}_{sg}\}_{i=1}^{4}$. Since the spatial resolutions of the semantic feature $f_{s}$ and the depth feature $f_{d}^{i}$ differ, we align $f_{s}$ to match $f_{d}^{i}$ by applying bilinear interpolation followed by max pooling, resulting in the resolution-aligned semantic feature $\tilde{f}_{s}$. The same resolution-aligned semantic feature $\tilde{f}_{s}$ is shared across all Bridging Gates. The decoder then leverages these semantic-aware geometric features to regress the depth map $I^{\prime}_{d}$ of the input image as follows:
$\begin{gathered}I^{\prime}_{d}=D_{d}(\mathbf{F}_{sg}),~{}\mathbf{F}_{sg}=\{G^{i}(f^{i}_{d},\tilde{f}_{s})\}_{i=1}^{4},\\ ~{}I^{\prime}_{d}\in\mathbb{R}^{H^{\prime}\times W^{\prime}},~{}F^{i}_{sg}\in\mathbb{R}^{HW\times C}.\end{gathered}$ (3)
From the following section, we simplify the notation by omitting the subscript $i$ (e.g., $f^{i}_{d}$ is written as $f_{d}$) to streamline the equations and discussion.
Figure 2: Overview of the proposed BriGeS framework. We design a pipeline that adaptively integrates semantic information into a depth foundation model through the Bridging Gate. During inference, this integration is further refined by Attention Temperature Scaling.
4.2 Bridging Gate
While DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] demonstrates superior generalization capabilities across a wide range of unseen scenes, its lack of semantic integration can result in over-smoothed predictions in scenarios involving complex structures or ambiguous boundaries. To address this, we design the Bridging Gate that enables direct interaction between geometric and semantic information. This enhancement aims to provide more accurate depth estimations, particularly in scenes characterized by complex structures or unclear object boundaries. Our Bridging Gate ($G$) consists of a cross-attention block ($\operatorname{Block}_{c}$) and a self-attention block ($\operatorname{Block}_{s}$), designed to dynamically fuse geometric and semantic information as follows:
$G(f_{d},\tilde{f}_{s})=\operatorname{Block}_{s}\left(\operatorname{Block}_{c}(f_{d},\tilde{f}_{s})\right).$ (4)
Cross-Attention Block. In the cross-attention block $\operatorname{Block}_{c}$, the depth feature $f_{d}$ is used as the query, while the semantic feature $\tilde{f}_{s}$ serves as both the key and value. This configuration allows for the direct fusion of geometric and semantic information. The resulting fused feature is further refined through an MLP layer to generate $F_{c}$ as follows:
$\begin{gathered}F_{c}=\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right),~{}\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{d}K_{s}^{T}}{\sqrt{d}}\right)\cdot V_{s}\right),\\ Q_{d}=W_{q}^{c}(f_{d}),~{}K_{s}=W_{k}^{c}(\tilde{f}_{s}),~{}V_{s}=W_{v}^{c}(\tilde{f}_{s}),\end{gathered}$ (5)
where $d$ denotes the dimension of the projected query. The weight matrices $W_{q}^{c}$, $W_{k}^{c}$, and $W_{v}^{c}$ are used to transform features $f_{d}$ and $\tilde{f}_{s}$ to the query, key, and value, respectively.
Self-Attention Block. Subsequently, the generated $F_{c}$ is further processed in the self-attention block $\operatorname{Block}_{s}$. In this step, $F_{c}$ serves as the query, key, and value, enabling the self-attention mechanism to refine the feature. The output is then passed through an MLP to produce the final semantic-aware geometric feature $F_{sg}$ as follows:
$\begin{gathered}F_{sg}=\operatorname{Block_{s}}\left(F_{c}\right),~{}\operatorname{Block_{s}}\left(F_{c}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{c}K_{c}^{T}}{\sqrt{d}}\right)\cdot V_{c}\right),\\ Q_{c}=W_{q}^{s}(F_{c}),\quad K_{c}=W_{k}^{s}(F_{c}),\quad V_{c}=W_{v}^{s}(F_{c}),\end{gathered}$ (6)
where $W_{q}^{s}$, $W_{k}^{s}$, and $W_{v}^{s}$ are the weight matrices that transforms a feature $F_{c}$ to the query, key and value, respectively. This feature represents a more precise fusion of geometric and semantic information.
4.3 Attention Temperature Scaling
The Bridging Gate effectively merges features from two different features ($f_{d}$, $\tilde{f}_{s}$) through attention mechanisms. However, it can neglect peripheral details within the $f_{d}$ feature, as attention may become overly concentrated on central object regions. To address this issue, we leverage the threshold $\tau$ within the Bridging Gate’s attention operations during the inference phase as follows:
$\begin{gathered}\operatorname{Attn}_{\tau}\left(Q,K,V\right)=\operatorname{softmax}\left(\frac{{QK^{T}}}{\tau\sqrt{d}}\right)\cdot V,~{}\tau>1,\end{gathered}$ (7)
where $Q$, $K$, and $V$ represent query, key, and value from either the cross-attention block ($Q_{d},K_{s},V_{s}$) or the self-attention block ($Q_{c},K_{c},V_{c}$). By dividing the scaled-dot product between the query and key by $\tau$, this approach effectively broadens the distribution of attention from highly concentrated areas to more peripheral regions, facilitating a more comprehensive assessment of both central and surrounding features. This redistribution of attention, applied during the inference phase, enhances the model’s ability to incorporate overall structure, thereby improving the overall depth estimation performance.
4.4 Training
Building a foundation model with high generalization capabilities typically requires extensive amounts of data and substantial computational resources. However, to improve performance with minimal data and effort, we adopt a strategy focused on optimizing key components effectively. Specifically, we train only the Bridging Gate, keeping the encoders ($E_{s}$, $E_{d}$) and decoder ($D_{d}$) parameters frozen, as shown in Fig. 2. This approach is designed to generate enhanced scene understanding features by combining geometric and semantic information within the Bridging Gate.
5 Experiments
5.1 Implementation Details
Our training methodology aligns closely with the one employed in the DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. Initially, the ground truth depth value is transformed into a disparity space and then normalized to a range of 0 to 1. To effectively train on datasets with varying depths, we employ an affine-invariant loss, which ignores the effects of scale and shift. When training BriGeS with DepthAnything-V1, only the affine-invariant loss is used, whereas training with DepthAnything-V2 incorporates both the affine-invariant loss and the gradient matching loss from MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] at a 1:2 ratio.
The batch size is set to 16, with an initial learning rate of $55e-$ 5. We use the AdamW optimizer [Loshchilov(2017)] with a linear schedule. We conduct experiments on 8 RTX TITAN GPUs using labeled images, running for 2 epochs with DepthAnything-V1 and 1 epoch with DepthAnything-V2, respectively. Training our model for a single epoch takes 8 hours for DepthAnything-Base and 18 hours for DepthAnything-Large. During training, all images are resized so that the shorter side is 518 while maintaining the original aspect ratio, and then randomly cropped to a size of 518$\times$518. In addition, the temperature scaling factor $\tau$ is set to 2.5 during the inference. Our model contains 244M parameters based on the DepthAnything-Base, and 744M based on the DepthAnything-Large, including the SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] encoder.
5.2 Experimental Setup
Training Datasets. For training BriGeS, we employ a carefully curated subset of the extensive datasets featured in DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. To enhance the model’s generalization capabilities, we select three diverse datasets encompassing both indoor and outdoor environments for each model. Our training data comprises 600K samples from BlendedMVS [Yao et al.(2020)Yao, Luo, Li, Zhang, Ren, Zhou, Fang, and Quan], HRWSI [Xian et al.(2020)Xian, Zhang, Wang, Mai, Lin, and Cao], and TartanAir [Wang et al.(2020c)Wang, Zhu, Wang, Hu, Qiu, Wang, Hu, Kapoor, and Scherer] for DepthAnything-V1, and 582K samples from VKITTI [Cabon et al.(2020)Cabon, Murray, and Humenberger], Hypersim [Roberts et al.(2021)Roberts, Ramapuram, Ranjan, Kumar, Bautista, Paczan, Webb, and Susskind], and TartanAir for DepthAnything-V2, which corresponds to approximately 1% of the training datasets originally used to train the respective DepthAnything.
Method Backbone KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus] ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger] DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]
AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$
MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] ResNeXt-101 0.236 0.630 0.111 0.885 0.184 0.752 0.332 0.715
DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun] ViT-Large 0.100 0.901 0.098 0.903 0.078 0.946 0.182 0.758
Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] Stable Diffusion 0.099 0.916 0.055 0.964 0.065 0.960 0.308 0.773
GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen] Stable Diffusion 0.080 0.934 0.058 0.969 0.096 0.959 0.226 0.741
DepthAnything-V1-Base [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] ViT-Base 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
DepthAnything-V1-Base + Ours ViT-Base 0.074 0.947 0.042 0.982 0.051 0.980 0.243 0.756
DepthAnything-V1-Large [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] ViT-Large 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
DepthAnything-V1-Large + Ours ViT-Large 0.076 0.950 0.040 0.984 0.050 0.983 0.222 0.761
DepthAnything-V2-Base [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] ViT-Base 0.084 0.936 0.046 0.977 0.062 0.980 0.281 0.75
DepthAnything-V2-Base + Ours ViT-Base 0.080 0.939 0.046 0.977 0.057 0.981 0.259 0.752
DepthAnything-V2-Large [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] ViT-Large 0.080 0.944 0.043 0.979 0.053 0.983 0.265 0.754
DepthAnything-V2-Large + Ours ViT-Large 0.075 0.948 0.044 0.979 0.049 0.985 0.249 0.754
Table 1: Quantitative comparison of zero-shot relative depth estimation. We compare with the state-of-the-art models that estimate affine-invariant depth. Our evaluation protocol aligns with Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler]. (Bold: best, Underlined: second best)
Table 2: Quantitative comparison of zero-shot depth estimation. Performance on DA-2K [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] benchmark. All metrics are accuracy (%). DA indicates DepthAnything. (Bold: best)
Evaluation Details. We perform zero-shot relative depth evaluation on four datasets: KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], and DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]. Since our method predicts affine-invariant inverse depth, we restrict comparison to relative depth estimation models in this evaluation to ensure a fair comparison. Evaluation is performed by aligning the estimated depth $\mathbf{d}^{*}$ with the ground truth $\mathbf{d}$ by computing scale and shift values. These values are then used to derive an aligned depth map $\tilde{\mathbf{d}}~{}(=\text{scale}\times\mathbf{d}^{*}+\text{shift})$, which is adjusted to the same unit as the ground truth. For quantitative assessment, we use the aligned depth and the ground truth, employing metrics such as AbsRel (absolute relative error: $\frac{1}{M}\sum_{i=1}^{M}\left|\tilde{\mathbf{d}}_{i}-\mathbf{d}_{i}\right|/\mathbf{d}_{i}$) and $\delta_{1}$ accuracy (percentage of pixels where $\max\left(\tilde{\mathbf{d}}_{i}/\mathbf{d}_{i},\mathbf{d}_{i}/\tilde{\mathbf{d}}_{i}\right)<1.25$). We further evaluate our method on DA-2K [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], a high-resolution benchmark comprising diverse scenes, which is well-suited for analyzing fine-grained details and generalization performance. In this benchmark, the accuracy is measured by identifying the closer of two distinct pixels. Given this evaluation protocol based on pairwise depth comparison, we extend our evaluation to include metric depth estimation models, enabling a comprehensive assessment of our method.
5.3 Comparison to State-of-the-art Methods
We evaluate our BriGeS on zero-shot depth estimation through both quantitative and qualitative comparisons. Table 1 provides a comparison of our model with six state-of-the-art relative depth estimation methods including MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun], DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun], Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler], GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen], DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] and V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] across multiple benchmark datasets [Geiger et al.(2012)Geiger, Lenz, and Urtasun, Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus, Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger, Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]. For the DA-2K benchmark [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], we extend the evaluation to include nine representative methods, covering both relative and metric depth estimation approaches. In addition to six models above, we include UniDepth [Piccinelli et al.(2024)Piccinelli, Yang, Sakaridis, Segu, Li, Van Gool, and Yu], Metric3d-V2 [Hu et al.(2024)Hu, Yin, Zhang, Cai, Long, Chen, Wang, Yu, Shen, and Shen], and DepthPro [Bochkovskii et al.(2025)Bochkovskii, Delaunoy, Germain, Santos, Zhou, Richter, and Koltun], as detailed in Table 2.
Figure 3: Qualitative results on unseen datasets of BriGeS (Ours) and various state-of-the-art methods. BriGeS is highly effective at capturing intricate structures and delicate objects, such as thin wires, tree branches, and nets. Best viewed when zoomed in.
Quantitative Comparison. As shown in Table 1, BriGeS consistently achieves enhanced performance across the majority of datasets. Compared to DepthAnything-V1 and V2, it achieves an average reduction of 7.33% in the AbsRel metric. Notably, the most significant performance improvement is observed in the DIODE dataset [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.], with an AbsRel reduction of 15.33% in the BriGeS + DepthAnything-V1-Base setup, highlighting its remarkable effectiveness. Similarly, in Table 2, BriGeS achieves improved performance on the DA-2K benchmark.
Temperature Scaling Factor 2 2.5 3 3.5 4
Avg. Rank (↓) 2.3 1.3 2 1.5 2.1
Table 3: Ablation study on the scaling factor $\tau$ in our Attention Temperature Scaling. (Bold: best)
[…]
figureQualitative results of ablation study on proposed modules. A reversed color map is applied for better visualization.
When built upon DepthAnything-V2, it achieves the highest performance on the DA-2K, surpassing other state-of-the-art methods, demonstrating that integrating geometric and semantic information significantly enhances depth estimation accuracy.
Qualitative Comparison. Fig. 1, 3 show representative depth estimation results from our method. These collectively indicate that existing models generally fail to capture semantic information for accurate depth estimation. The first row from Fig. 3 showcases the inference results of thin power lines, which are best preserved in our method, demonstrating its capacity to recover delicate structures. Similarly, in the second row, our model captures the intricate and thin structure of tree branches more effectively than DepthAnything-V2, while in the third row, it best preserves the fine-grained details of the fishing net. Although Marigold and GenPercept, which leverage Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], perform well with fine details, they mispredict the sky, as seen in the first and second rows. These results show that our BriGeS effectively fuses geometric and semantic information, accurately delineating object edges at similar depths and preserving fine details in complex structures.
5.4 Ablation Study
We conduct ablation studies on BriGeS using both DepthAnything-V1-Base and Large [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] to assess the effectiveness and robustness of our proposed methods across different scales.
The quantitative results are presented in Table 4 for the base and large models, showcasing consistent performance improvements. As evident in Table 4-(b), our Bridging Gate effectively compensates for the lack of semantic information in the geometric features across different model sizes. This improvement is consistently observed in both models, demonstrating the versatility and effectiveness of our Bridging Gate in enhancing semantic details irrespective of the model’s scale.
Model size # Component KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus] ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger] DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]
Bridging Gate Temperature Scaling AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑
(a) 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
Base (b) 0.078 0.944 0.043 0.982 0.054 0.979 0.246 0.760
(c) 0.074 0.947 0.042 0.982 0.051 0.980 0.243 0.756
(a) 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
Large (b) 0.076 0.950 0.042 0.984 0.054 0.983 0.226 0.760
(c) 0.076 0.950 0.040 0.984 0.050 0.983 0.222 0.761
Table 4: Ablation study on proposed modules. (a) The baseline model DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao]. (b) The baseline with Bridging Gate. (c) Our method includes the Bridging Gate and the Attention Temperature Scaling. (Bold: best, underlined: second best)
Additionally, Table 4-(c) demonstrates that our Attention Temperature Scaling effectively enhances the model’s overall performance in both model sizes. These results confirm that the highest performance is achieved when all proposed modules are applied, demonstrating the synergistic effect of our approaches.
Table 3 presents the average rank calculated for the quantitative results across KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], and DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.] by varying the temperature scaling factor for BriGeS based on DepthAnything-V1-Large. We empirically set the scaling factor to 2.5, as it yields the best performance. The qualitative results in Fig. 3 further underscore the impact of our methods. Notably, without the scaling module, the tower cranes within the black box were not well predicted, an issue that is effectively resolved with the scaling module. These results demonstrate that the Bridging Gate and the scaling module help distribute attention more effectively. By mitigating excessive focus on central objects, they reduce prediction errors associated with small structures, thereby enhancing overall performance.
6 Conclusion
In this work, we introduce Bridging Geometric and Semantic (BriGeS), an effective approach designed to enhance Monocular Depth Estimation (MDE) by integrating geometric and semantic information through foundation models. The core of our method, the Bridging Gate, utilizes cross-attention and self-attention mechanisms to merge depth and semantic data effectively. Additionally, our Attention Temperature Scaling technique ensures a balanced distribution of attention across varied inputs. By leveraging pre-trained models, BriGeS requires minimal training data, thereby achieving both resource efficiency and strong robustness across diverse scenarios. Our comprehensive evaluations demonstrate that BriGeS significantly outperforms existing methods, particularly in complex scenes, thus, we establish a new standard for the MDE foundation model by effectively combining geometric and semantic insights. However, its reliance on two foundation models entails a trade-off in memory efficiency, which we aim to address by distilling knowledge into an integrated encoder that directly produces a semantic-aware geometric representation in the future.
Supplementary Material
Appendix A More Qualitative Results
For a detailed examination of the qualitative performance across four unseen datasets, please refer to Fig. 4 through Fig. 7. These figures present the results for the KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.], respectively. In this section, we compare our model with DepthAnything-V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. Our model outperforms particularly in complex scenes, such as those containing intricate structures and thin objects.
Figure 4: Qualitative results on KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun].
Figure 5: Qualitative results on NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus].
Figure 6: Qualitative results on ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger].
Figure 7: Qualitative results on DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.].

### v2

@@
-Engineering and Computer Science, DGIST, South Korea Bridging Geometric and Semantic Foundation Models
+
 Abstract
@@
 Extensive experiments across multiple challenging datasets demonstrate that BriGeS outperforms state-of-the-art methods in MDE for complex scenes, effectively handling intricate structures and overlapping objects.
-1 Introduction
-Figure 1: Qualitative comparison with DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao]. Our method excels in capturing more complex visual elements and details.
+I Introduction
+Figure 1: Qualitative comparison with DepthAnything-V1 [23].
 Monocular Depth Estimation (MDE) is a crucial task that predicts the depth map from a single image.
@@
 This adaptability makes it especially valuable in fields like robotics, autonomous driving, and augmented reality.
-Historically, MDE relies on handcrafted features and traditional vision techniques [Hoiem et al.(2007)Hoiem, Efros, and Hebert, Liu et al.(2008)Liu, Yuen, Torralba, Sivic, and Freeman, Saxena et al.(2008)Saxena, Sun, and Ng].
-However, the advent of CNN [Eigen et al.(2014)Eigen, Puhrsch, and Fergus, Laina et al.(2016)Laina, Rupprecht, Belagiannis, Tombari, and Navab, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Wang et al.(2020b)Wang, Zhang, Wang, Lu, and Ruan, Ricci et al.(2018)Ricci, Ouyang, Wang, Sebe, et al., Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow] and ViT-based models [Yang et al.(2021)Yang, Tang, Ding, Sebe, and Ricci, Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun, Yuan et al.(2022)Yuan, Gu, Dai, Zhu, and Tan, Zhao et al.(2022)Zhao, Zhang, Poggi, Tosi, Guo, Zhu, Huang, Tang, and Mattoccia, Cheng et al.(2021)Cheng, Zhang, and Tang, Bae et al.(2023)Bae, Moon, and Im] has significantly enhanced its performance.
-Recent studies [Kendall et al.(2017)Kendall, Martirosyan, Dasgupta, Henry, Kennedy, Bachrach, and Bry, Klingner et al.(2020)Klingner, Termöhlen, Mikolajczyk, and Fingscheidt, Wang et al.(2020a)Wang, Zhang, Wang, Lin, and Lu, Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li, Xing et al.(2023)Xing, Shen, Ho, and Tzes, Shvets et al.(2024)Shvets, Zhao, Niethammer, Sengupta, and Berg, Li et al.(2023)Li, Xue, Su, He, Mao, Zhu, Sun, and Zhang, Choi et al.(2020)Choi, Jung, Lee, and Kim, Sanchez-Escobedo et al.(2018)Sanchez-Escobedo, Lin, Casas, and Pardas, Zama Ramirez et al.(2019)Zama Ramirez, Poggi, Tosi, Mattoccia, and Di Stefano, Jiao et al.(2018)Jiao, Cao, Song, and Lau, Chen et al.(2019)Chen, Liu, Liu, and Wang] have highlighted the advantages of incorporating semantic context from images, which provide more informative cues than solely relying on local geometry.
-More recently, MDE works [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have focused on training large-scale models using a vast amount of training datasets inspired by foundation models [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al., Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun, Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.].
+Historically, MDE relies on handcrafted features and traditional vision techniques [6].
+However, the advent of CNN [4] and ViT-based models [13] has significantly enhanced its performance.
+Recent studies [10] have highlighted the advantages of incorporating semantic context from images, which provide more informative cues than solely relying on local geometry.
+More recently, MDE works [23, 24, 8] have focused on training large-scale models using a vast amount of training datasets inspired by foundation models [2, 14, 9].
 These approaches aim to harness the broader contextual understanding inherent in large pre-trained models, significantly advancing the field of depth estimation.
-Despite these advancements, current foundation models in MDE [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have primarily utilized geometric data for depth estimation and have not directly incorporated semantic information.
+Despite these advancements, current foundation models in MDE [23, 24, 8] have primarily utilized geometric data for depth estimation and have not directly incorporated semantic information.
 These models require substantial training data and considerable computational resources, making it challenging to integrate additional semantic information.
@@
 1.
-To integrate depth features with semantic features under a resource-constrained training environment, we design Bridging Gate, which directly fuses features from the depth [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] and segmentation [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] foundation model.
+To integrate depth features with semantic features under a resource-constrained training environment, we design Bridging Gate, which directly fuses features from the depth [23, 24] and segmentation [9] foundation model.
 This gate employs cross-attention and self-attention blocks to facilitate effective feature integration.
@@
   • We introduce the Attention Temperature Scaling, which regulates the attention distribution to mitigate over-concentration on specific regions during inference.
-2 Related Works
-Monocular Depth Estimation.
-Monocular Depth Estimation (MDE) method estimates depth from a single image by learning the spatial shape and structure of objects.
-Some pioneering research in this field predominantly utilizes handcrafted features and traditional vision techniques [Hoiem et al.(2007)Hoiem, Efros, and Hebert, Liu et al.(2008)Liu, Yuen, Torralba, Sivic, and Freeman, Saxena et al.(2008)Saxena, Sun, and Ng].
-However, these methods often rely solely on local information and face challenges in capturing comprehensive scene structure.
-To address these challenges, many MDE models have adopted CNN [Eigen et al.(2014)Eigen, Puhrsch, and Fergus, Laina et al.(2016)Laina, Rupprecht, Belagiannis, Tombari, and Navab, Ricci et al.(2018)Ricci, Ouyang, Wang, Sebe, et al., Wang et al.(2020b)Wang, Zhang, Wang, Lu, and Ruan] and ViT-based approaches [Yang et al.(2021)Yang, Tang, Ding, Sebe, and Ricci, Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun, Yuan et al.(2022)Yuan, Gu, Dai, Zhu, and Tan, Bae et al.(2023)Bae, Moon, and Im].
-Recently, research has pivoted towards a self-supervised approach [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Zhou et al.(2022)Zhou, Taylor, Greenwood, and Mackiewicz, Fan et al.(2023)Fan, Tosi, Poggi, and Mattoccia, Saunders et al.(2024)Saunders, Manso, and Vogiatzis] due to the high cost of labeling training data, allowing training on larger datasets.
-While these models face challenges with objects having unclear boundaries or similar depths [Klingner et al.(2020)Klingner, Termöhlen, Mikolajczyk, and Fingscheidt, Dinh and Lim(2023), Shvets et al.(2024)Shvets, Zhao, Niethammer, Sengupta, and Berg], numerous studies have started integrating semantic information to refine edge predictions [Wang et al.(2020a)Wang, Zhang, Wang, Lin, and Lu, Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li, Xing et al.(2023)Xing, Shen, Ho, and Tzes].
-For instance, ROIFormer [Xing et al.(2023)Xing, Shen, Ho, and Tzes] proposes a semantic-aware region of interest transformer that uses geometric cues from semantic segmentation to guide the attention mechanism.
-TriDepth [Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li] further introduces a semantic-aware triplet loss to address the edge-fattening problem, which significantly improves model performance while incurring no additional computational cost.
-Zero-shot Depth Estimation.
-Zero-shot depth estimation involves predicting depth from images unseen during training.
-Earlier efforts in this domain focused on collecting more training images to enhance the generalization capabilities of MDE models [Chen et al.(2016)Chen, Fu, Yang, and Deng, Xian et al.(2018)Xian, Shen, Cao, Lu, Xiao, Li, and Luo, Chen et al.(2020)Chen, Qian, Fan, Kojima, Hamilton, and Deng].
-However, data variability in unconstrained settings often leads to inconsistent depth prediction due to differing depth scales and shifts.
-To resolve these problems, several methods are designed to estimate affine-invariant depth, which adjusts for global scale and shift discrepancies without knowing the precise object size.
-A pioneering work, MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] proposes an affine-invariant loss to facilitate effective zero-shot transfer across diverse datasets by estimating relative depth.
-DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun] also adapts the same loss term, and further enhances the performance of zero-shot MDE by replacing CNN-based architecture with vision transformers [Dosovitskiy(2020)].
-For metric depth estimation, ZoeDepth [Bhat et al.(2023)Bhat, Birkl, Wofk, Wonka, and Müller] integrates relative and metric depth supervision, and Metric3D [Yin et al.(2023)Yin, Zhang, Chen, Cai, Yu, Wang, Chen, and Shen, Hu et al.(2024)Hu, Yin, Zhang, Cai, Long, Chen, Wang, Yu, Shen, and Shen] introduces a canonical camera space transformation module to resolve metric ambiguity.
-Recent works leverage the rich visual knowledge from the pre-trained model to perform zero-shot knowledge transfer.
-For instance, DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] develops a semi-supervised training strategy using DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] encoder, while Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] and GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen] use a model derived from Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer].
-3 Preliminary: Foundation Models
-In this section, we provide a brief overview of the foundation model used in our pipeline.
-DepthAnything.
-The depth foundation model, DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] employs a ViT-based encoder $E_{d}$ and a CNN-based decoder $D_{d}$.
-The encoder processes the input image $I$ to generate features $\{f^{i}_{d}\}_{i=1}^{4}$, which serve as the input for the decoder.
-The decoder then utilizes this feature to regress the depth map $I_{d}$ of the input image as follows:
-$\begin{gathered}I_{d}=D_{d}(\{f^{i}_{d}\}_{i=1}^{4}),~{}\{f^{i}_{d}\}_{i=1}^{4}=E_{d}(I),\\ I_{d}\in\mathbb{R}^{H^{\prime}\times W^{\prime}},~{}f^{i}_{d}\in\mathbb{R}^{HW\times C},~{}I\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3}\end{gathered}$ (1)
-where $H$, $W$, and $C$ denote the height, width, and channel of the feature map, respectively, while $H^{\prime}$ and $W^{\prime}$ represent the height and width of the input image and the resulting depth.
-SegmentAnything.
-The segmentation foundation model, SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] consists of a ViT-based encoder $E_{s}$, a prompt encoder $E_{p}$, and a decoder $D_{s}$.
-The encoder $E_{s}$ processes the input image $I$ to generate a feature $f_{s}$.
-The prompt encoder takes an input $x$, which can be points, boxes, or text, and generates a corresponding prompt token $p$.
-Finally, the decoder combines the feature map $f_{s}$ with the prompt token $p$ to generate the mask $M$ for the input image as follows:
-$\begin{gathered}M=D_{s}(f_{s},p),~{}f_{s}=E_{s}(I),~{}p=E_{p}(x),\\ ~{}M\in\mathbb{R}^{H^{\prime}\times W^{\prime}},~{}f_{s}\in\mathbb{R}^{C\times H\times W},I\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3},\end{gathered}$ (2)
-where $H^{\prime}$ and $W^{\prime}$ represent the height and width of masks.
-4 Method
-4.1 Overall Pipeline
+II Method
+II-A Overall Pipeline
 In this section, we describe the overall pipeline of BriGeS, illustrated in Fig.
 2.
-Our BriGeS utilizes the encoders $E_{d}$ from DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] and $E_{s}$ from SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.], along with the DepthAnything decoder $D_{d}$.
+Our BriGeS utilizes the encoders $E_{d}$ from DepthAnything [23, 24] and $E_{s}$ from SegmentAnything [9], along with the DepthAnything decoder $D_{d}$.
 We extract depth features $\{f^{i}_{d}\}_{i=1}^{4}$ and semantic feature $f_{s}$ from the respective encoders, and fuse them using Bridging Gates $\{G^{i}\}_{i=1}^{4}$ to generate semantic-aware geometric features $\mathbf{F}_{sg}=\{F^{i}_{sg}\}_{i=1}^{4}$.
@@
 From the following section, we simplify the notation by omitting the subscript $i$ (e.g., $f^{i}_{d}$ is written as $f_{d}$) to streamline the equations and discussion.
-Figure 2: Overview of the proposed BriGeS framework.
-We design a pipeline that adaptively integrates semantic information into a depth foundation model through the Bridging Gate.
-During inference, this integration is further refined by Attention Temperature Scaling.
-4.2 Bridging Gate
-While DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] demonstrates superior generalization capabilities across a wide range of unseen scenes, its lack of semantic integration can result in over-smoothed predictions in scenarios involving complex structures or ambiguous boundaries.
+II-B Bridging Gate
+While DepthAnything [23, 24] demonstrates superior generalization capabilities across a wide range of unseen scenes, its lack of semantic integration can result in over-smoothed predictions in scenarios involving complex structures or ambiguous boundaries.
 To address this, we design the Bridging Gate that enables direct interaction between geometric and semantic information.
@@
 The resulting fused feature is further refined through an MLP layer to generate $F_{c}$ as follows:
-$\begin{gathered}F_{c}=\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right),~{}\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{d}K_{s}^{T}}{\sqrt{d}}\right)\cdot V_{s}\right),\\ Q_{d}=W_{q}^{c}(f_{d}),~{}K_{s}=W_{k}^{c}(\tilde{f}_{s}),~{}V_{s}=W_{v}^{c}(\tilde{f}_{s}),\end{gathered}$ (5)
+$\begin{gathered}F_{c}=\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right),\\ \operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{d}K_{s}^{T}}{\sqrt{d}}\right)\cdot V_{s}\right),\\ Q_{d}=W_{q}^{c}(f_{d}),~K_{s}=W_{k}^{c}(\tilde{f}_{s}),~V_{s}=W_{v}^{c}(\tilde{f}_{s}),\end{gathered}$ (3)
 where $d$ denotes the dimension of the projected query.
@@
 The output is then passed through an MLP to produce the final semantic-aware geometric feature $F_{sg}$ as follows:
-$\begin{gathered}F_{sg}=\operatorname{Block_{s}}\left(F_{c}\right),~{}\operatorname{Block_{s}}\left(F_{c}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{c}K_{c}^{T}}{\sqrt{d}}\right)\cdot V_{c}\right),\\ Q_{c}=W_{q}^{s}(F_{c}),\quad K_{c}=W_{k}^{s}(F_{c}),\quad V_{c}=W_{v}^{s}(F_{c}),\end{gathered}$ (6)
+$\begin{gathered}F_{sg}=\operatorname{Block_{s}}\left(F_{c}\right),\\ \operatorname{Block_{s}}\left(F_{c}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{c}K_{c}^{T}}{\sqrt{d}}\right)\cdot V_{c}\right),\\ Q_{c}=W_{q}^{s}(F_{c}),\quad K_{c}=W_{k}^{s}(F_{c}),\quad V_{c}=W_{v}^{s}(F_{c}),\end{gathered}$ (4)
 where $W_{q}^{s}$, $W_{k}^{s}$, and $W_{v}^{s}$ are the weight matrices that transforms a feature $F_{c}$ to the query, key and value, respectively.
 This feature represents a more precise fusion of geometric and semantic information.
-4.3 Attention Temperature Scaling
+Figure 2: Overview of the proposed BriGeS framework.
+II-C Attention Temperature Scaling
 The Bridging Gate effectively merges features from two different features ($f_{d}$, $\tilde{f}_{s}$) through attention mechanisms.
@@
 This redistribution of attention, applied during the inference phase, enhances the model’s ability to incorporate overall structure, thereby improving the overall depth estimation performance.
-4.4 Training
+II-D Training
 Building a foundation model with high generalization capabilities typically requires extensive amounts of data and substantial computational resources.
@@
 This approach is designed to generate enhanced scene understanding features by combining geometric and semantic information within the Bridging Gate.
-5 Experiments
-5.1 Implementation Details
-Our training methodology aligns closely with the one employed in the DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao].
+III Experiments
+TABLE I: Quantitative comparison of zero-shot relative depth estimation. (Bold: best, Underlined: second best)
+Method Backbone KITTI NYUv2 ETH3D DIODE
+AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$
+MiDaS ResNeXt-101 0.236 0.630 0.111 0.885 0.184 0.752 0.332 0.715
+DPT ViT-Large 0.100 0.901 0.098 0.903 0.078 0.946 0.182 0.758
+Marigold Stable Diffusion 0.099 0.916 0.055 0.964 0.065 0.960 0.308 0.773
+GenPercept Stable Diffusion 0.080 0.934 0.058 0.969 0.096 0.959 0.226 0.741
+DepthAnything-V1-Base ViT-Base 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
+DepthAnything-V1-Base + Ours ViT-Base 0.074 0.947 0.042 0.982 0.051 0.980 0.243 0.756
+DepthAnything-V1-Large ViT-Large 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
+DepthAnything-V1-Large + Ours ViT-Large 0.076 0.950 0.040 0.984 0.050 0.983 0.222 0.761
+DepthAnything-V2-Base ViT-Base 0.084 0.936 0.046 0.977 0.062 0.980 0.281 0.750
+DepthAnything-V2-Base + Ours ViT-Base 0.080 0.939 0.046 0.977 0.057 0.981 0.259 0.752
+DepthAnything-V2-Large ViT-Large 0.080 0.944 0.043 0.979 0.053 0.983 0.265 0.754
+DepthAnything-V2-Large + Ours ViT-Large 0.075 0.948 0.044 0.979 0.049 0.985 0.249 0.754
+TABLE II: Comparison of zero-shot depth estimation on DA-2K [24]. All metrics are accuracy (%). (Bold: best, DA: DepthAnything)
+Relative depth estimation model
+MiDaS [14] DPT [13] Marigold [8] GenPercept [22] DA-V1-B [23] DA-V1-L [23] DA-V2-B [24] DA-V2-L [24]
+83.5 79.1 86.8 93.2 89.4 88.5 97.0 97.1
+Metric depth estimation model BriGeS (Ours)
+UniDepth [12] Metric3D-V2 [7] DepthPro [1] DA-V1-B DA-V1-L DA-V2-B DA-V2-L
+83.7 89.2 94.7 91.4 92.1 97.0 97.6
+III-A Implementation Details
+Our training methodology aligns closely with the one employed in the DepthAnything [23, 24].
 Initially, the ground truth depth value is transformed into a disparity space and then normalized to a range of 0 to 1.
 To effectively train on datasets with varying depths, we employ an affine-invariant loss, which ignores the effects of scale and shift.
-When training BriGeS with DepthAnything-V1, only the affine-invariant loss is used, whereas training with DepthAnything-V2 incorporates both the affine-invariant loss and the gradient matching loss from MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] at a 1:2 ratio.
-The batch size is set to 16, with an initial learning rate of $55e-$ 5.
-We use the AdamW optimizer [Loshchilov(2017)] with a linear schedule.
+When training BriGeS with DepthAnything-V1, only the affine-invariant loss is used, whereas training with DepthAnything-V2 incorporates both the affine-invariant loss and the gradient matching loss from MiDaS [14] at a 1:2 ratio.
+The batch size is set to 16, with an initial learning rate of 5e55e-5.
+We use the AdamW optimizer [11] with a linear schedule.
 We conduct experiments on 8 RTX TITAN GPUs using labeled images, running for 2 epochs with DepthAnything-V1 and 1 epoch with DepthAnything-V2, respectively.
 Training our model for a single epoch takes 8 hours for DepthAnything-Base and 18 hours for DepthAnything-Large.
-During training, all images are resized so that the shorter side is 518 while maintaining the original aspect ratio, and then randomly cropped to a size of 518$\times$518.
+During training, all images are resized so that the shorter side is 518 while maintaining the original aspect ratio, and then randomly cropped to a size of $\times518$.
 In addition, the temperature scaling factor $\tau$ is set to 2.5 during the inference.
-Our model contains 244M parameters based on the DepthAnything-Base, and 744M based on the DepthAnything-Large, including the SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] encoder.
-5.2 Experimental Setup
+Our model contains 244M parameters based on the DepthAnything-Base, and 744M based on the DepthAnything-Large, including the SegmentAnything [9] encoder.
+III-B Experimental Setup
 Training Datasets.
-For training BriGeS, we employ a carefully curated subset of the extensive datasets featured in DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao].
+For training BriGeS, we employ a carefully curated subset of the extensive datasets featured in DepthAnything [23, 24].
 To enhance the model’s generalization capabilities, we select three diverse datasets encompassing both indoor and outdoor environments for each model.
-Our training data comprises 600K samples from BlendedMVS [Yao et al.(2020)Yao, Luo, Li, Zhang, Ren, Zhou, Fang, and Quan], HRWSI [Xian et al.(2020)Xian, Zhang, Wang, Mai, Lin, and Cao], and TartanAir [Wang et al.(2020c)Wang, Zhu, Wang, Hu, Qiu, Wang, Hu, Kapoor, and Scherer] for DepthAnything-V1, and 582K samples from VKITTI [Cabon et al.(2020)Cabon, Murray, and Humenberger], Hypersim [Roberts et al.(2021)Roberts, Ramapuram, Ranjan, Kumar, Bautista, Paczan, Webb, and Susskind], and TartanAir for DepthAnything-V2, which corresponds to approximately 1% of the training datasets originally used to train the respective DepthAnything.
-Method Backbone KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus] ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger] DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]
-AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$
-MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] ResNeXt-101 0.236 0.630 0.111 0.885 0.184 0.752 0.332 0.715
-DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun] ViT-Large 0.100 0.901 0.098 0.903 0.078 0.946 0.182 0.758
-Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] Stable Diffusion 0.099 0.916 0.055 0.964 0.065 0.960 0.308 0.773
-GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen] Stable Diffusion 0.080 0.934 0.058 0.969 0.096 0.959 0.226 0.741
-DepthAnything-V1-Base [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] ViT-Base 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
-DepthAnything-V1-Base + Ours ViT-Base 0.074 0.947 0.042 0.982 0.051 0.980 0.243 0.756
-DepthAnything-V1-Large [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] ViT-Large 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
-DepthAnything-V1-Large + Ours ViT-Large 0.076 0.950 0.040 0.984 0.050 0.983 0.222 0.761
-DepthAnything-V2-Base [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] ViT-Base 0.084 0.936 0.046 0.977 0.062 0.980 0.281 0.75
-DepthAnything-V2-Base + Ours ViT-Base 0.080 0.939 0.046 0.977 0.057 0.981 0.259 0.752
-DepthAnything-V2-Large [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] ViT-Large 0.080 0.944 0.043 0.979 0.053 0.983 0.265 0.754
-DepthAnything-V2-Large + Ours ViT-Large 0.075 0.948 0.044 0.979 0.049 0.985 0.249 0.754
-Table 1: Quantitative comparison of zero-shot relative depth estimation.
-We compare with the state-of-the-art models that estimate affine-invariant depth.
-Our evaluation protocol aligns with Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler].
-(Bold: best, Underlined: second best)
-Table 2: Quantitative comparison of zero-shot depth estimation.
-Performance on DA-2K [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] benchmark.
-All metrics are accuracy (%).
-DA indicates DepthAnything.
-(Bold: best)
+Our training data comprises 600K samples from BlendedMVS [25], HRWSI [21], and TartanAir [20] for DepthAnything-V1, and 582K samples from VKITTI [3], Hypersim [15], and TartanAir for DepthAnything-V2, which corresponds to approximately 1% of the training datasets originally used to train the respective DepthAnything.
 Evaluation Details.
-We perform zero-shot relative depth evaluation on four datasets: KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], and DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.].
+We perform zero-shot relative depth evaluation on four datasets: KITTI [5], NYUv2 [18], ETH3D [17], and DIODE [19].
 Since our method predicts affine-invariant inverse depth, we restrict comparison to relative depth estimation models in this evaluation to ensure a fair comparison.
@@
 For quantitative assessment, we use the aligned depth and the ground truth, employing metrics such as AbsRel (absolute relative error: $\frac{1}{M}\sum_{i=1}^{M}\left|\tilde{\mathbf{d}}_{i}-\mathbf{d}_{i}\right|/\mathbf{d}_{i}$) and $\delta_{1}$ accuracy (percentage of pixels where $\max\left(\tilde{\mathbf{d}}_{i}/\mathbf{d}_{i},\mathbf{d}_{i}/\tilde{\mathbf{d}}_{i}\right)<1.25$).
-We further evaluate our method on DA-2K [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], a high-resolution benchmark comprising diverse scenes, which is well-suited for analyzing fine-grained details and generalization performance.
+We further evaluate our method on DA-2K [24], a high-resolution benchmark comprising diverse scenes, which is well-suited for analyzing fine-grained details and generalization performance.
 In this benchmark, the accuracy is measured by identifying the closer of two distinct pixels.
 Given this evaluation protocol based on pairwise depth comparison, we extend our evaluation to include metric depth estimation models, enabling a comprehensive assessment of our method.
-5.3 Comparison to State-of-the-art Methods
+Figure 3: Qualitative results on unseen datasets of BriGeS (Ours) and various state-of-the-art methods.
+TABLE III: Ablation study on proposed modules.
+(a) The baseline model DepthAnything-V1 [23].
+(b) The baseline with Bridging Gate.
+(c) Our method includes the Bridging Gate and the Attention Temperature Scaling.
+(Bold: best, underlined: second best)
+Model size # Component KITTI [5] NYUv2 [18] ETH3D [17] DIODE [19]
+Bridging Gate Temperature Scaling AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑
+(a) 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
+Base (b) 0.078 0.944 0.043 0.982 0.054 0.979 0.246 0.760
+(c) 0.074 0.947 0.042 0.982 0.051 0.980 0.243 0.756
+(a) 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
+Large (b) 0.076 0.950 0.042 0.984 0.054 0.983 0.226 0.760
+(c) 0.076 0.950 0.040 0.984 0.050 0.983 0.222 0.761
+III-C Comparison to State-of-the-art Methods
 We evaluate our BriGeS on zero-shot depth estimation through both quantitative and qualitative comparisons.
-Table 1 provides a comparison of our model with six state-of-the-art relative depth estimation methods including MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun], DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun], Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler], GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen], DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] and V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] across multiple benchmark datasets [Geiger et al.(2012)Geiger, Lenz, and Urtasun, Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus, Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger, Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.].
-For the DA-2K benchmark [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], we extend the evaluation to include nine representative methods, covering both relative and metric depth estimation approaches.
-In addition to six models above, we include UniDepth [Piccinelli et al.(2024)Piccinelli, Yang, Sakaridis, Segu, Li, Van Gool, and Yu], Metric3d-V2 [Hu et al.(2024)Hu, Yin, Zhang, Cai, Long, Chen, Wang, Yu, Shen, and Shen], and DepthPro [Bochkovskii et al.(2025)Bochkovskii, Delaunoy, Germain, Santos, Zhou, Richter, and Koltun], as detailed in Table 2.
-Figure 3: Qualitative results on unseen datasets of BriGeS (Ours) and various state-of-the-art methods.
-BriGeS is highly effective at capturing intricate structures and delicate objects, such as thin wires, tree branches, and nets.
-Best viewed when zoomed in.
+Table I provides a comparison of our model with six state-of-the-art relative depth estimation methods including MiDaS [14], DPT [13], Marigold [8], GenPercept [22], DepthAnything-V1 [23] and V2 [24] across multiple benchmark datasets [5, 18, 17, 19].
+For the DA-2K benchmark [24], we extend the evaluation to include nine representative methods, covering both relative and metric depth estimation approaches.
+In addition to six models above, we include UniDepth [12], Metric3d-V2 [7], and DepthPro [1], as detailed in Table II.
 Quantitative Comparison.
-As shown in Table 1, BriGeS consistently achieves enhanced performance across the majority of datasets.
+As shown in Table I, BriGeS consistently achieves enhanced performance across the majority of datasets.
 Compared to DepthAnything-V1 and V2, it achieves an average reduction of 7.33% in the AbsRel metric.
-Notably, the most significant performance improvement is observed in the DIODE dataset [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.], with an AbsRel reduction of 15.33% in the BriGeS + DepthAnything-V1-Base setup, highlighting its remarkable effectiveness.
-Similarly, in Table 2, BriGeS achieves improved performance on the DA-2K benchmark.
+Notably, the most significant performance improvement is observed in the DIODE dataset [19], with an AbsRel reduction of 15.33% in the BriGeS + DepthAnything-V1-Base setup, highlighting its remarkable effectiveness.
+Similarly, in Table II, BriGeS achieves improved performance on the DA-2K benchmark.
+When built upon DepthAnything-V2, it achieves the highest performance on the DA-2K, surpassing other state-of-the-art methods, demonstrating that integrating geometric and semantic information significantly enhances depth estimation accuracy.
+TABLE IV: Ablation study on the scaling factor $\tau$ in our Attention Temperature Scaling. (Bold: best)
 Temperature Scaling Factor 2 2.5 3 3.5 4
 Avg. Rank (↓) 2.3 1.3 2 1.5 2.1
-Table 3: Ablation study on the scaling factor $\tau$ in our Attention Temperature Scaling. (Bold: best)
-[…]
-figureQualitative results of ablation study on proposed modules. A reversed color map is applied for better visualization.
-When built upon DepthAnything-V2, it achieves the highest performance on the DA-2K, surpassing other state-of-the-art methods, demonstrating that integrating geometric and semantic information significantly enhances depth estimation accuracy.
+Figure 4: Qualitative results of ablation study on proposed modules.
 Qualitative Comparison.
@@
 Similarly, in the second row, our model captures the intricate and thin structure of tree branches more effectively than DepthAnything-V2, while in the third row, it best preserves the fine-grained details of the fishing net.
-Although Marigold and GenPercept, which leverage Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], perform well with fine details, they mispredict the sky, as seen in the first and second rows.
+Although Marigold and GenPercept, which leverage Stable Diffusion [16], perform well with fine details, they mispredict the sky, as seen in the first and second rows.
 These results show that our BriGeS effectively fuses geometric and semantic information, accurately delineating object edges at similar depths and preserving fine details in complex structures.
-5.4 Ablation Study
-We conduct ablation studies on BriGeS using both DepthAnything-V1-Base and Large [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] to assess the effectiveness and robustness of our proposed methods across different scales.
-The quantitative results are presented in Table 4 for the base and large models, showcasing consistent performance improvements.
-As evident in Table 4-(b), our Bridging Gate effectively compensates for the lack of semantic information in the geometric features across different model sizes.
+III-D Ablation Study
+We conduct ablation studies on BriGeS using both DepthAnything-V1-Base and Large [23] to assess the effectiveness and robustness of our proposed methods across different scales.
+The quantitative results are presented in Table III for the base and large models, showcasing consistent performance improvements.
+As evident in Table III-(b), our Bridging Gate effectively compensates for the lack of semantic information in the geometric features across different model sizes.
 This improvement is consistently observed in both models, demonstrating the versatility and effectiveness of our Bridging Gate in enhancing semantic details irrespective of the model’s scale.
-Model size # Component KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus] ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger] DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]
-Bridging Gate Temperature Scaling AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑ AbsRel↓ $\delta$1↑
-(a) 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
-Base (b) 0.078 0.944 0.043 0.982 0.054 0.979 0.246 0.760
-(c) 0.074 0.947 0.042 0.982 0.051 0.980 0.243 0.756
-(a) 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
-Large (b) 0.076 0.950 0.042 0.984 0.054 0.983 0.226 0.760
-(c) 0.076 0.950 0.040 0.984 0.050 0.983 0.222 0.761
-Table 4: Ablation study on proposed modules.
-(a) The baseline model DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao].
-(b) The baseline with Bridging Gate.
-(c) Our method includes the Bridging Gate and the Attention Temperature Scaling.
-(Bold: best, underlined: second best)
-Additionally, Table 4-(c) demonstrates that our Attention Temperature Scaling effectively enhances the model’s overall performance in both model sizes.
+Additionally, Table III-(c) demonstrates that our Attention Temperature Scaling effectively enhances the model’s overall performance in both model sizes.
 These results confirm that the highest performance is achieved when all proposed modules are applied, demonstrating the synergistic effect of our approaches.
-Table 3 presents the average rank calculated for the quantitative results across KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], and DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.] by varying the temperature scaling factor for BriGeS based on DepthAnything-V1-Large.
+Table IV presents the average rank calculated for the quantitative results across KITTI [5], NYUv2 [18], ETH3D [17], and DIODE [19] by varying the temperature scaling factor for BriGeS based on DepthAnything-V1-Large.
 We empirically set the scaling factor to 2.5, as it yields the best performance.
@@
 By mitigating excessive focus on central objects, they reduce prediction errors associated with small structures, thereby enhancing overall performance.
-6 Conclusion
+IV Conclusion
 In this work, we introduce Bridging Geometric and Semantic (BriGeS), an effective approach designed to enhance Monocular Depth Estimation (MDE) by integrating geometric and semantic information through foundation models.
@@
 However, its reliance on two foundation models entails a trade-off in memory efficiency, which we aim to address by distilling knowledge into an integrated encoder that directly produces a semantic-aware geometric representation in the future.
-Supplementary Material
-Appendix A More Qualitative Results
-For a detailed examination of the qualitative performance across four unseen datasets, please refer to Fig.
-4 through Fig.
-7.
-These figures present the results for the KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.], respectively.
-In this section, we compare our model with DepthAnything-V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao].
-Our model outperforms particularly in complex scenes, such as those containing intricate structures and thin objects.
-Figure 4: Qualitative results on KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun].
-Figure 5: Qualitative results on NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus].
-Figure 6: Qualitative results on ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger].
-Figure 7: Qualitative results on DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.].